OpenStudy (anonymous):

The decomposition of SO2Cl2 is a first order process with a half-life of 5000 seconds. If a reaction starts with 0.750 M SO2Cl2, how long will it take for the concentration to decrease to 0.500 M?

OpenStudy (aaronq):

use the first order law equation, the integrated form is the easiest. \(ln[A]=-kt+ln[A]_0\)

OpenStudy (aaronq):

you can find the rate constant, k, with \(t_{1/2}=\dfrac{ln(2)}{k}\)

OpenStudy (anonymous):

thank you!

OpenStudy (aaronq):

no problem !